OpenStudy (anonymous):

which of the following is the equation of a line parallel to the line y=-x+1, passing through the point (4,1) ? Please help ? A. x+y=-5 B. x-y=-5 C. x+y=5 D. -x-y=5

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

the equation is y=mx+b and you would fill in the numbers

OpenStudy (anonymous):

I'm still pretty confused , I know that formula but i've never done a problem like this , is there anyway you could give more information

OpenStudy (anonymous):

oh wait i have no idea!! i thought i did this is different hmm i just tried and got it wrong.. im sorry i really dont know try looking through other questions similar and tagging the people that helped them

OpenStudy (anonymous):

It's okay thank you

OpenStudy (anonymous):

y=-x+1 what do you need to do to both sides?

OpenStudy (anonymous):

I'm not really sure , I'm really confused

OpenStudy (anonymous):

y=-x+1 +x +x y+x=1

OpenStudy (anonymous):

Ohh thank you so much ! So the answer is C. ?

OpenStudy (anonymous):

One second the 5 is throwing me off

OpenStudy (anonymous):

yeah it is :P

OpenStudy (anonymous):

thank you !!

OpenStudy (anonymous):

mmhm no problem!